# Machine Learning INDI Incremental Nonlinear Dynamic Inversion

### v2

@@ Abstract
This paper presents a method to recover quadrotor Unmanned Air Vehicles from a throw, when no control parameters are known before the throw. We leverage the availability of high-frequency rotor speed feedback available in racing drone hardware and software to find control effectiveness values and fit a motor model using recursive least squares (RLS) estimation. Furthermore, we propose an excitation sequence that provides large actuation commands while guaranteeing to stay within gyroscope sensing limits. After 450ms of excitation, an Incremental Nonlinear Dynamic Inversion (INDI) attitude controller uses the 52 fitted parameters to arrest rotational motion and recover an upright attitude. Finally, a Nonlinear Dynamic Inversion (NDI) position controller drives the craft to a position setpoint. The proposed algorithm runs efficiently on microcontrollers found in common UAV flight controllers, and was shown to recover an agile quadrotor every time in 57 live experiments with as low as $3.5$m throw height, demonstrating robustness against initial rotations and noise. We also demonstrate control of randomized quadrotors in simulated throws, where the parameter fitting Root-Mean-Square (RMS) error is typically within 10% of the true value.
@@ I INTRODUCTION
Figure 1: Composite image: The quadrotor learns to fly before it would hit the ground.
@@ II-A Controller
where $\nu_{0}$ is a measurement of $\nu$ from onboard sensors.
@@ II-B Model Identification
Each motor is successively excited with two step inputs followed by a decreasing ramp. This ensures that many different levels of $\delta$, and the largest possible $\dot{\omega}$ are provided to the system. On entry to the first step of each motor, the gyroscope reading $\Omega_{e}$ and margin $\Delta\Omega$ to the sensor limits is recorded. The excitation is aborted if a gyroscope reading exceeds $\Omega_{e}+\Delta\Omega/n_{l}$ where $n_{l}$ is the number of motors left to excite after the current motor. It could be argued that this is too conservative, since if a quadrotor is able to hover, then 2 of its motors must provide a rotation that de-saturates the sensor, so $0.5n_{l}$ in the denominator should be sufficient. However, if we assume that rotor acceleration and deceleration have the same dynamics, then in the worst case, the thrust from a rotor imparts as much angular impulse on spin-up as it does while spinning down after the excitation is aborted. This effectively halves the permissible margin $\Delta\Omega$.
@@
II-C Simulation Setup
The open-source robot simulator Gazebo is used to obtain parameterizable quadrotor simulations. The motor locations, time constant, moment constant and input non-linearity $\kappa$ were drawn from uniform distributions, while ensuring that the resulting vehicle is still physically capable of hover.
A plugin based on the Open Software Robotics Foundation’s vehicle_gateway project111github.com/osrf/vehicle_gateway was then used to run 16 parameter-randomized simulations with the flight controller in the loop. The initial conditions were a parabolic trajectory of approximately $4$m height. $450$ms after the launch, the excitation sequence is triggered, and after that the described controller drives the vehicle to a point $1.5$m above the origin.
@@ II-C Simulation Setup
II-D Experimental Setup
The UAV was instrumented with reflective markers to enable optical position tracking. The positioning data was sent to the UAV, but only used for position control, not identification. Aside from the positioning triangulation and transmitting that data to the drone, all calculations are performed onboard. The internal attitude estimation remains the Mahony filter provided by Betaflight. Heading, velocity and position use simple constant gain observers to fuse inertial measurements with optical position and velocity data. If less accurate or reliable positioning systems like GNSS are to be used for future outdoor experiments, these estimators would likely have to be replaced by an Extended Kalman Filter.
Conditions
The nominal INDI controller parameters were calculated with bench tests of propeller and motors, and measurements of geometry and mass properties. For controlled experiments, this controller took the quadrotor from ground level onto a $3.5$-$4$m high parabolic trajectory and induced a pseudo-random rotation between 0 and 10rad/s. Using the known geometry of the quadrotor, and after performing bench tests with a controller-motor-propeller combination, the parameters were calculated. The engines are then spooled down and 250ms later the excitation is triggered. A further 450ms later, the fitting is complete and switchover to the identified parameter-set occurs.
When manually throwing, the release is detected using the accelerometer, and a longer delay time of 500ms is used.
III RESULTS
III-A Simulations
All $16$ vehicles were successfully stabilized despite errors in the parameter estimates. Table II shows that the root-mean-square (RMS) errors between the true and estimated parameters over the $16$ runs are typically below $10$%. The motor numbers $1$ to $4$ correspond to the order in which they have been excited. For reference, the “nominal” column provides the mean of the absolute value of the $16$ realizations of the true simulation parameters.
TABLE II: Parameter error RMS over the $16$ simulated runs.
@@ III-A Simulations
$B_{1,x}\cdot k\cdot 10^{6}$ 0.000 0.008 0.003 0.008 0.005
$B_{1,y}\cdot k\cdot 10^{6}$ 0.000 0.005 0.006 0.009 0.002
$B_{1,z}\cdot k\cdot 10^{6}$ 0.621 0.022 0.034 0.037 0.044
$B_{1,p}\cdot k\cdot 10^{6}$ 22.455 2.441 2.001 2.191 1.714
$B_{1,q}\cdot k\cdot 10^{6}$ 21.912 2.145 1.267 1.422 1.224
$B_{1,r}\cdot k\cdot 10^{6}$ 4.770 0.224 0.441 0.527 0.204
$B_{2,p}\cdot 10^{3}$ 0.000 0.055 0.092 0.102 0.031
$B_{2,q}\cdot 10^{3}$ 0.000 0.066 0.075 0.077 0.023
$B_{2,r}\cdot 10^{3}$ 1.011 0.070 0.068 0.077 0.057
$\omega_{\text{max}}$ 4113 193.3 103.9 141.3 141.4
$\kappa$ 0.470 0.064 0.147 0.232 0.184
$\omega_{\text{idle}}$ 0 7.141 15.84 30.32 14.88
$\tau$ [ms] 23.835 4.793 4.797 8.335 5.227
Figure 5: Angular rate tracking and motor commands during recovery
III-B Experiments
A total of 57 controlled launches were performed, and during all of them the parameters were identified well enough to stabilize the UAVs attitude.
Figure 4 shows the time evolution of the 8 parameters in the model for $\Delta\dot{r}$ during an arbitrarily chosen experiment run (the 4 parameters in the center plot relate to each motor velocity and 4 parameters in the right plot relate to motor acceleration). The low-pass filtering is effective in reducing noise in the measured targets shown in the left plot. From the center and right plot it can be seen that for all motors but motor 1 the parameters initially do not converge. This is due to the presence of only idle rotation speed before the step input. However, once excitation is provided, the parameters immediately converge, and reproduction of the rotation acceleration measurement is unaffected by the initial uncertainty. Nonetheless, control based on these parameters would be impossible.
@@ III-B Experiments
Finally, table I shows the means and standard deviations of the fitted parameters over the last 20 runs with identical conditions. They are compared to the calculated “reference” value from bench tests, geometry and mass properties measurements. Note, that these calculations also have their own uncertainty and should not be seen as a ground truth. In particular, it is reasonable that motors 1 and 3 (the rear motors) have higher yaw effectiveness $B_{1,r}$ and higher thrust effectiveness $B_{1,z}$ than motors 2 and 4, because the structure that the motors are mounted on is much wider in the front of our quadrotor. This can reduce the downwards flow velocity and swirl.

### v3

@@ Abstract
This paper presents a method to recover quadrotor Unmanned Air Vehicles from a throw, when no control parameters are known before the throw. We leverage the availability of high-frequency rotor speed feedback available in racing drone hardware and software to find control effectiveness values and fit a motor model using recursive least squares (RLS) estimation. Furthermore, we propose an excitation sequence that provides large actuation commands while guaranteeing to stay within gyroscope sensing limits. After 450ms of excitation, an Incremental Nonlinear Dynamic Inversion (INDI) attitude controller uses the 52 fitted parameters to arrest rotational motion and recover an upright attitude. Finally, a Nonlinear Dynamic Inversion (NDI) position controller drives the craft to a position setpoint. The proposed algorithm runs efficiently on microcontrollers found in common UAV flight controllers, and was shown to recover an agile quadrotor every time in live experiments with as low as $3.5$m throw height, demonstrating robustness against initial rotations and noise. We also demonstrate control of randomized quadrotors in simulated throws, where the parameter fitting Root-Mean-Square (RMS) error is typically within 10% of the true value.
@@ I INTRODUCTION
Figure 1: Composite image: The quadrotor learns to fly before it would hit the ground. Moving picture at https://youtu.be/CLFPXcpzA14
@@ II-A Controller
where $\nu_{0}$ is a measurement of $\nu$ from onboard sensors. $B_{1}$ and $B_{2}$ are called “control effectiveness” matrices and arise from the geometry and actuator parameters of the vehicle.
@@ II-B Model Identification
Each motor is successively excited with two step inputs followed by a decreasing ramp. This ensures that many different levels of $\delta$, and the largest possible $\dot{\omega}$ are provided to the system. On entry to the first step of each motor, the gyroscope reading $\Omega_{e}$ and margin $\Delta\Omega$ to the sensor limits is recorded. The excitation is aborted if a gyroscope reading exceeds $\Omega_{e}+\Delta\Omega/n_{l}$ where $n_{l}$ is the number of motors left to excite after the current motor. This is illustrated in figure 3. It could be argued that this is too conservative, since if a quadrotor is able to hover, then 2 of its motors must provide a rotation that de-saturates the sensor, so $0.5n_{l}$ in the denominator should be sufficient. However, if we assume that rotor acceleration and deceleration have the same dynamics, then in the worst case, the thrust from a rotor imparts as much angular impulse on spin-up as it does while spinning down after the excitation is aborted. This effectively halves the permissible margin $\Delta\Omega$.
@@
Figure 5: Angular rate tracking and motor commands during recovery
II-C Experimental Setup
The UAV was instrumented with reflective markers to enable optical position tracking. The positioning data was sent to the UAV, but only used for position control; it does not influence identification. Aside from the positioning triangulation and transmitting that data to the drone, all calculations are performed onboard. The internal attitude estimation remains the Mahony filter provided by Betaflight. Heading, velocity and position use simple constant gain observers to fuse inertial measurements with optical position and velocity data.
Conditions
The nominal INDI controller parameters were calculated with bench tests of propeller and motors, and measurements of geometry and mass properties. For controlled experiments, this controller took the quadrotor from ground level onto a $3.5$-$4$m high parabolic trajectory and induced a pseudo-random rotation between 0 and 10rad/s. Using the known geometry of the quadrotor, and after performing bench tests with a controller-motor-propeller combination, the parameters were calculated. The engines are then spooled down for 250ms, and the excitation is triggered. After its completion 450ms later, the fitting is done and switchover to the identified parameter-set occurs.
II-D Simulation Setup
[…]
The simulation is linked against the same INDIflight software as used in the experiments, and can be reproduced using Docker333github.com/tudelft/indiflightSupport/tree/iros_imav_2024. All data files are also available44410.4121/0530be90-cc6c-4029-9774-670657882906.
III RESULTS
@@
TABLE II: Parameter error RMS over the $1000$ simulated runs.
@@
$B_{1,x}\cdot k\cdot 10^{6}$ 0.000 0.037 0.036 0.032 0.037
$B_{1,y}\cdot k\cdot 10^{6}$ 0.000 0.038 0.034 0.032 0.036
$B_{1,z}\cdot k\cdot 10^{6}$ 0.463 0.040 0.036 0.033 0.039
$B_{1,p}\cdot k\cdot 10^{6}$ 18.301 1.147 1.515 1.597 1.379
$B_{1,q}\cdot k\cdot 10^{6}$ 17.538 1.107 1.518 1.535 1.424
$B_{1,r}\cdot k\cdot 10^{6}$ 4.173 0.472 0.436 0.395 0.395
$B_{2,p}\cdot 10^{3}$ 0.000 0.055 0.068 0.085 0.114
$B_{2,q}\cdot 10^{3}$ 0.000 0.055 0.064 0.089 0.116
$B_{2,r}\cdot 10^{3}$ 0.667 0.030 0.032 0.039 0.052
$\omega_{\text{max}}$ 4900 390.7 275.7 148.7 47.46
$\kappa$ 0.499 0.119 0.105 0.114 0.101
$\omega_{\text{idle}}$ 0.000 2.824 5.568 11.28 4.660
$\tau$ [ms] 25.04 0.627 0.689 0.768 1.080
III-A Experiments
A total of 57 controlled launches were performed, and during all of them the parameters were identified well enough to stabilize the UAVs attitude. Due to an oversight, suitable log files were only obtained in the last 20 runs.
Figure 4 shows the time evolution of the 8 parameters in the model for $\Delta\dot{r}$ during an experiment run (the 4 parameters in the center plot relate to each motor velocity and 4 parameters in the right plot relate to motor acceleration). The low-pass filtering is effective in reducing noise in the measured targets shown in the left plot. From the center and right plot it can be seen that for all motors but motor 1 the parameters initially do not converge. This is due to the presence of only idle rotation speed before the step input. However, once excitation is provided, the parameters immediately converge, and reproduction of the rotation acceleration measurement is unaffected by the initial uncertainty. Nonetheless, control based on these parameters would be impossible.
@@ III-A Experiments
Finally, table I shows the means and standard deviations of the fitted parameters over the last 20 runs. They are compared to the calculated “reference” value from bench tests, geometry and mass properties measurements. Note, that these calculations also have their own uncertainty and should not be seen as a ground truth. In particular, it is reasonable that motors 1 and 3 (the rear motors) have higher yaw effectiveness $B_{1,r}$ and higher thrust effectiveness $B_{1,z}$ than motors 2 and 4, because the structure that the motors are mounted on is much wider in the front of our quadrotor. This can reduce the downwards flow velocity and swirl.
@@
III-B Simulations
All $1000$ vehicles were successfully stabilized despite errors in the parameter estimates. Table II shows that the root-mean-square (RMS) errors between the true and estimated parameters over the runs are typically below $10$%. The motor numbers $1$ to $4$ correspond to the order in which they have been excited. The “nominal” column gives context to interpret the RMS values, and is the mean of the absolute parameter value over all the true generated crafts.